OpenStudy (anonymous):

You volunteer to help drive children at a charity event to the zoo, but you can fit only 6 of the 15 children present in your van. How many different groups of 6 children can you drive?

jimthompson5910 (jim_thompson5910):

use the formula n C r = (n!)/(r!*(n-r)!) where in this case, n = 15 and r = 6

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

sweet thanks

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

i got 5005

jimthompson5910 (jim_thompson5910):

you are 100% correct

OpenStudy (anonymous):

sweet!! next question

OpenStudy (anonymous):

There are 6 performers who will present their comedy acts this weekend at a comedy club. One of the performers insists on being the last stand-up comic of the evening. If this performer's request is granted, how many different ways are there to schedule the appearances?

jimthompson5910 (jim_thompson5910):

one insists to be last, so that slot is taken you have 6-1 = 5 slots left to be filled with 5 performers

jimthompson5910 (jim_thompson5910):

order matters so there are 5 P 5 = (5!)/((5-5)!) = ??? ways

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

120?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

yes!

jimthompson5910 (jim_thompson5910):

an alternative is to compute 5! = 5*4*3*2*1 = 120

OpenStudy (anonymous):

im finally getting this

jimthompson5910 (jim_thompson5910):

that's great

jimthompson5910 (jim_thompson5910):

glad it's all clicking now

OpenStudy (anonymous):

ive been working on this for the past month. thank you so much!!

jimthompson5910 (jim_thompson5910):

hmm maybe it just takes a month for things to settle into place?

jimthompson5910 (jim_thompson5910):

but glad to be of help

OpenStudy (anonymous):

haha thank you! im taking an online class so i have teach myself everything. its been crazy

jimthompson5910 (jim_thompson5910):

I'm sure you have a teacher to go to though right?

OpenStudy (anonymous):

yeah.. but i cant actually go see here. im a visual learner...

jimthompson5910 (jim_thompson5910):

like if you have any questions

jimthompson5910 (jim_thompson5910):

lol who isn't really, I know I am

OpenStudy (anonymous):

exactly

jimthompson5910 (jim_thompson5910):

you can at least contact the teacher though right?